# FedX: Explanation-Guided Pruning for Communication-Efficient Federated Learning in Remote Sensing

### v1

@@ V Experimental Results
We carried out different kinds of experiments in order to: 1) analyze FedX on different pruning configurations and relevance-based explanation methods; 2) evaluate FedX across different DL architectures; 3) assess the communication and computational efficiency of FedX across different DL architectures; and 4) compare FedX with the state-of-the-art FL algorithms.
V-A Analysis of FedX with Different Pruning Configurations and Relevance-Based Explanation Methods
In this subsection, we assess the performance of FedX, considering different pruning configurations and relevance-based explanation methods. To this end, we exploit ResNet18 as the backbone model due to its moderate number of parameters. As mentioned in Section IV-B, we exploit LRP [30] and IG [31] as explanation methods for the experiments, considering layer-wise and global pruning configurations. This results in four different scenarios: i) FedX with LRP and the layer-wise configuration, denoted as FedX (LRP, Layerwise); ii) FedX with LRP and the global configuration, denoted as FedX (LRP, Global); iii) FedX with IG and the layer-wise configuration, denoted as FedX (IG, Layerwise); iv) FedX with IG and the global configuration, denoted as FedX (IG, Global). Fig. 2 shows the obtained mAP scores for the BigEarthNet-S2 dataset.
@@ V-A Analysis of FedX with Different Pruning Configurations and Relevance-Based Explanation Methods
Figure 2: mAP scores obtained by FedX with different pruning configurations (layer-wise and global) and relevance-based explanation methods (LRP and IG) for ResNet18 under varying pruning rates (the BigEarthNet-S2 dataset).
@@
Figure 4: mAP scores obtained by FedX (using the LRP method with a layer-wise configuration) for different model architectures under varying pruning rates for: (a) the BigEarthNet-S2; and (b) the EuroSAT datasets.
From Fig. 2, one can also observe that FedX (IG, Layerwise) and FedX (LRP, Layerwise) provide similar mAP scores when the pruning rate is between 10% and 40%. Within this range, the performance remains relatively stable, with mAP values fluctuating between 63% and 67%. FedX (IG, Layerwise) performs slightly better than FedX (LRP, Layerwise) at pruning rates of %30 and 50%, reaching 66% and 70% respectively. However, FedX (LRP, Layerwise) outperforms FedX (IG, Layerwise) at pruning rates of 10% and 60%. As an example, when the pruning rate is set to 60%, FedX (LRP, Layerwise) achieves an mAP score of 69%, while FedX (IG, Layerwise) reaches 63%. Once the pruning rate exceeds 70%, the performance of FedX (IG, Layerwise) begins to degrade more rapidly than that of FedX (LRP, Layerwise). From this point onward, FedX (LRP, Layerwise) shows better robustness across the higher pruning range. At the highest pruning rate of 90%, FedX (LRP, Layerwise) still retains 55% mAP, whereas FedX (IG, Layerwise) drops further to 53%. One can observe that despite the decline, FedX (LRP, Layerwise) retains a higher proportion of the original performance under extreme sparsity conditions. This behavior highlights the comparative robustness of FedX (LRP, Layerwise) in preserving relevant model structure when the number of model parameters is significantly reduced. While FedX (IG, Layerwise) shows better performance at certain pruning rates, FedX (LRP, Layerwise) performs better overall. Based on its superior robustness and consistent performance, we use FedX (LRP, Layerwise) for the rest of the experiments, denoting it as FedX for simplicity.
@@ V-D Comparison of FedX with State-of-the-Art Algorithms
In this subsection, we compare FedX with random pruning and two state-of-the-art algorithms: i) FedDST [28] and ii) FedDIP [29]. The comparison is carried out on different pruning rates using the ResNet6 architecture. This architecture was chosen due to its relatively small number of parameters, which limits its capacity to compensate for pruning through overparameterization. This choice also ensures a fairer comparison of pruning strategies and allows a clearer understanding of their impact on recovery and convergence under high sparsity. Fig. 5 shows the corresponding mAP scores for the BigEarthNet-S2 dataset.
From the results, one can see that FedX consistently yields higher mAP values than random pruning and both state-of-the-art algorithms across almost all sparsity levels. As an example, at a pruning rate of 20%, FedX achieves an mAP score of 64%, outperforming the randomly pruned model by 6% and exceeding FedDST and FedDIP by approximately 4%. In greater detail, as the pruning rate increases from 20% to 50%, all the considered methods show a gradual decline in mAP. However, FedX achieves the highest performance throughout this range. At the 50% pruning rate, FedX achieves an mAP score of 63%, outperforming random pruning by 7%, FedDIP by 3%, and FedDST by 5%. Beyond this point, the performance gap widens with FedX, especially at extreme sparsity. As an example, at a pruning rate of 80%, FedX reaches 60% mAP, outperforming random pruning by 11% and both FedDST and FedDIP by around 5%. Even at a 90% pruning rate, FedX maintains competitive performance, while the comparison methods degrade substantially. This indicates that FedX is more effective than the state-of-the-art algorithms and random pruning in preserving essential model components, particularly in scenarios where model capacity is severely constrained. Across all pruning rates, FedX consistently maintains performance closer to the unpruned model compared to the random pruning and state-of-the-art algorithms. This shows that FedX can maintain performance remarkably close to the unpruned model, even at extreme sparsity rates. These results demonstrate the robustness of FedX in FL settings with constrained communication and model size, and highlight its ability to guide pruning in a performance-preserving manner. A similar behavior was also observed by using the EuroSAT dataset (not reported due to space constraints).
Figure 5: mAP scores obtained by FedX (using the LRP method with a layer-wise configuration), FedDST and FedDIP for ResNet6 under varying pruning rates (the BigEarthNet-S2 dataset).

### v2

@@ V Experimental Results
We carried out different kinds of experiments in order to: 1) analyze FedX on different pruning configurations and pruning methods; 2) evaluate FedX across different DL architectures; 3) assess the communication and computational efficiency of FedX across different DL architectures; and 4) compare FedX with the state-of-the-art FL algorithms.
V-A Analysis of FedX with Different Pruning Configurations and Pruning Methods
In this subsection, we assess the performance of FedX, considering different pruning configurations and pruning methods. To this end, we exploit ResNet18 as the backbone model due to its moderate number of parameters. As mentioned in Section IV-B, we exploit LRP [2] and IG [34] as explanation methods for the experiments, considering layer-wise and global pruning configurations. This results in five different scenarios: i) FedX with LRP and the layer-wise configuration, denoted as FedX (LRP, Layerwise); ii) FedX with LRP and the global configuration, denoted as FedX (LRP, Global); iii) FedX with IG and the layer-wise configuration, denoted as FedX (IG, Layerwise); iv) FedX with IG and the global configuration, denoted as FedX (IG, Global); and v) FedX with magnitude-based pruning, denoted as FedX (Magnitude), where components are pruned according to the L1 norm of the weights without relevance estimation. Fig. 2 shows the obtained mAP scores for the BigEarthNet-S2 dataset.
@@ V-A Analysis of FedX with Different Pruning Configurations and Pruning Methods
Figure 2: mAP scores obtained by FedX with different pruning configurations (layer-wise and global), relevance-based explanation methods (LRP and IG), and a magnitude-based pruning method for ResNet18 under varying pruning rates (the BigEarthNet-S2 dataset).
@@
Figure 4: Result scores obtained by FedX (using the LRP method with a layer-wise configuration) for different model architectures under varying pruning rates for: (a) the BigEarthNet-S2 (mAP); and (b) the EuroSAT (accuracy) datasets.
From Fig. 2, one can also observe that FedX (IG, Layerwise) and FedX (LRP, Layerwise) provide similar mAP scores when the pruning rate is between 10% and 40%. Within this range, the performance remains relatively stable, with mAP values fluctuating between 63% and 67%. FedX (IG, Layerwise) performs slightly better than FedX (LRP, Layerwise) at pruning rates of 30% and 50%, reaching 66% and 70% respectively. However, FedX (LRP, Layerwise) outperforms FedX (IG, Layerwise) at pruning rates of 10% and 60%. As an example, when the pruning rate is set to 60%, FedX (LRP, Layerwise) achieves an mAP score of 69%, while FedX (IG, Layerwise) reaches 63%. Once the pruning rate exceeds 70%, the performance of FedX (IG, Layerwise) begins to degrade more rapidly than that of FedX (LRP, Layerwise). From this point onward, FedX (LRP, Layerwise) shows better robustness across the higher pruning range. At the highest pruning rate of 90%, FedX (LRP, Layerwise) still retains 55% mAP, whereas FedX (IG, Layerwise) drops further to 53%. One can observe that despite the decline, FedX (LRP, Layerwise) retains a higher proportion of the original performance under extreme sparsity conditions. This behavior highlights the comparative robustness of FedX (LRP, Layerwise) in preserving relevant model structure when the number of model parameters is significantly reduced. While FedX (IG, Layerwise) shows better performance at certain pruning rates, FedX (LRP, Layerwise) performs better overall. In addition, FedX (Magnitude) performs comparably to the explanation-guided variants at low sparsity levels (up to 20% pruning). However, its performance drops sharply beyond 30% pruning and remains around 32% mAP for higher pruning rates. This highlights that selecting components solely based on weight magnitude is not sufficient to preserve task-relevant capacity under stronger pruning, and confirms the benefit of explanation-guided relevance estimation in FedX. Based on its superior robustness and consistent performance, we use FedX (LRP, Layerwise) for the rest of the experiments, denoting it as FedX for simplicity.
@@ V-D Comparison of FedX with State-of-the-Art Algorithms
In this subsection, we compare FedX with random pruning and four state-of-the-art algorithms: FedDST [3], FedDIP [27], LotteryFL [21], and PruneFL [17]. The comparison is carried out on different pruning rates using the ResNet6 architecture. This architecture was chosen due to its relatively small number of parameters, which limits its capacity to compensate for pruning through overparameterization. This choice also ensures a fairer comparison of pruning strategies and allows a clearer understanding of their impact on recovery and convergence under high sparsity. Fig. 5 shows the corresponding mAP scores for the BigEarthNet-S2 dataset.
From the results, one can see that FedX consistently yields higher mAP values than random pruning and the considered state-of-the-art algorithms across almost all sparsity levels. As an example, at a pruning rate of 20%, FedX achieves an mAP score of 64%, outperforming the randomly pruned model by 6% and exceeding FedDST, FedDIP, and PruneFL by approximately 4%. In greater detail, as the pruning rate increases from 20% to 50%, all the considered methods show a gradual decline in mAP. However, FedX achieves the highest performance throughout this range. At the 50% pruning rate, FedX achieves an mAP score of 63%, outperforming random pruning by 7%, FedDIP by 3%, FedDST by 5%, PruneFL by 10%, and LotteryFL by 20%. Beyond this point, the performance gap widens with FedX, especially at extreme sparsity. As an example, at a pruning rate of 80%, FedX reaches 60% mAP, outperforming random pruning by 11% and FedDST, FedDIP, and PruneFL by around 5%, while the performance gap compared to LotteryFL is substantially larger at this pruning level. Even at a 90% pruning rate, FedX maintains competitive performance, while the comparison methods degrade substantially. This indicates that FedX is more effective than the state-of-the-art algorithms and random pruning in preserving essential model components, particularly in scenarios where model capacity is severely constrained. Across all pruning rates, FedX consistently maintains performance closer to the unpruned model compared to the random pruning and state-of-the-art algorithms. This shows that FedX can maintain performance remarkably close to the unpruned model, even at extreme sparsity rates. These results demonstrate the robustness of FedX in FL settings with constrained communication and model size, and highlight its ability to guide pruning in a performance-preserving manner. A similar behavior was also observed by using the EuroSAT dataset (not reported due to space constraints).
Figure 5: mAP scores obtained by FedX (using the LRP method with a layer-wise configuration), FedDST, FedDIP, PruneFL, LotteryFL, and random pruning for ResNet6 under varying pruning rates (the BigEarthNet-S2 dataset).